A girl is pressing a 2-kg book against a vertical wall. If the coefficient of friction between the book and the wall is 0.30, how hard should the girl push to allow the book to slide down the wall at a constant speed?


1
Expert's answer
2021-02-08T18:37:55-0500

Let's apply the Newton's Second Law of Motion:


"\sum F_y=ma_y=0,""F_{fr}-mg=0,""\mu N-mg=0,""\mu F-mg=0,""F=\dfrac{mg}{\mu}=\dfrac{2\ kg\cdot9.8\ \dfrac{m}{s^2}}{0.30}=65.3\ N."